# What One Cannot, Two Can: Two-Layer Transformers Provably Represent Induction Heads on Any-Order Markov Chains

### v1

@@ 3 Warm Up: Construction with Two Heads in Layer One, One in Layer Two
As a warm up, in this section, we prove our first main result that we can represent $k^{\text{th}}$-order Markov processes with a $2$-layer, 22-head transformer. Towards the same, we recall that the $k^{\text{th}}$-order in-context counting estimator, i.e. conditional $k$-gram model, defined in Sec. 2.1 is closely related to the classical Laplacian smoothing, which in turn is the optimal Bayes estimator for the next-token predictive distribution Laplace (1814); Rissanen (1984). In view of this fact, in order to estimate the optimal conditional distribution of each token in-context, it is natural to ask: how does a transformer represent this conditional $k$-gram model? To this end, Rajaraman et al. (2024) introduces the notion of a $k$-th order induction head. This mechanism enables the attention layer of a transformer to assign at each time step the highest attention weight to past tokens whose length-$k$ context matches that of the current token. For the sake of completeness, we recall the formal definition.
@@
The salience of this higher-order induction head is immediately reflected by the fact that such a mechanism is indeed what is precisely needed to implement the conditional $k$-gram estimator from Def. 2.1. Capitalizing on this, Rajaraman et al. (2024) proves the best known result for higher-order Markov processes, through a $3$-layer, 11-head transformer architecture, which we recall below to motivate our main result:
Theorem 1. (Theorem 4 in Rajaraman et al. (2024)): The conditional $k$-gram model can be represented using a transformer with three layers, each containing a single attention head. The layers are separated by MLP blocks and include relative positional encodings as well as layer normalization. The embedding dimension scales as $\mathcal{O}(S)$.
@@
We provide a brief outline of the proof of Theorem 3. For each position $n$ in the sequence $s_{0:T}$, the first attention layer effectively learns a hard attention map concentrated on the indices $\{n,n-1,\ldots,n-k+1\}$. Specifically, within a subspace of its embedding space, this layer computes $\textbf{u}_{n}=\left({\sum_{i=0}^{k-1}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=0}^{k-1}3^{i}}\right)$, with attention weights $\operatorname{attn}_{n,i}=3^{i}/\left({\sum_{j=0}^{k-1}3^{j}}\right)$ for $i\in\{n,n-1,\ldots,n-k+1\}$, and zero otherwise. Intuitively, this $\textbf{u}_{n}$ allows to capture the past-$k$ context starting at position $n$. Similarly, the second attention layer focuses on the positions $\{n-1,n-2,\ldots,n-k\}$, and computes $\textbf{v}_{n}=\left({\sum_{i=1}^{k}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=1}^{k}3^{i}}\right)$, capturing the context shifted by one position. After these two layers, the embedding at position $n$ contains both $\textbf{v}_{n}/\|\textbf{v}_{n}\|_{2}$ and the normalized vector $\textbf{u}_{n}/\|\textbf{u}_{n}\|_{2}$, assuming appropriate choices of embedding dimensions, query/key/value matrices, positional encodings, and MLP layers with layer normalization. In particular, the architecture can be configured to compute the $\ell_{2}$-norm via layer norm, as detailed in App. A.1. The third attention layer then functions as a $k$-th order induction head: at position $n$, the attention score is made proportional to the cosine similarity ${\langle\textbf{u}_{T},\textbf{v}_{n}\rangle}/\left({\|\textbf{u}_{T}\|_{2}\|\textbf{v}_{n}\|_{2}}\right)$. This quantity equals 1 when $\textbf{u}_{T}=\textbf{v}_{n}$, and is strictly less than 1 otherwise. As the softmax temperature tends to infinity, or equivalently, as the attention logits are scaled by a constant tending to infinity, the model approaches the behavior of a conditional $k$-gram estimator. Hence, the third layer effectively implements a $k$-th order induction head. ∎
@@
(Depth-width tradeoff) Thm. 3 thus reveals an interesting tradeoff between depth (number of attention layers) and width (maximum number of attention heads per layer). While our construction here has both depth and width two, its counterpart in Thm. 3 instead has depth three and width one. Thus our new architecture trades off depth with an additional attention head, thereby increasing the width.
@@
We provide a brief sketch of the proof for Theorem 3, using the same notation as that of the Theorem 3, and focusing on the key differences between the two architectures.
@@ 4 Main Result: Two-Layer Single-Head Construction
Figure 2: Attention maps learnt by a two-layer, single-head transformer trained on sequences generated by random Markov chains of order 33 (Sec. 2.1). (i) in the first layer, the attention map shows a clear pattern: attention weights increase monotonically along the first three lower-diagonals and drop to zero beyond that. This suggests that the relative positional bias is maximized the diagonal with an offset of k=3-k=-3, i.e., the third diagonal below the main diagonal, which is consistent with the construction in Sec. 4, (ii) in the second layer, the attention map closely resembles the ideal attention pattern required to approximate the conditional $k$-gram estimator. We refer to App. H for further experimental details.
While Thm. 3 above demonstrates that the conditional $k$-gram can be represented by a two-layer transformer, it however relies on a two-head construction, thus still leaving open our motivating question: can a $2$-layer, 11-head transformer represent the conditional $k$-gram? In this section, we precisely address this gap and realize a $2$-layer, 11-head model to represent $k^{\text{th}}$-order Markov processes for any $k\geq 1$. The key idea behind our approach is to leverage the MLP—and specifically, the non-linearities such as ReLU and LayerNorm—to isolate symbols at particular positions that are useful for estimating induction heads. This contrasts with prior constructions, which primarily focused on the attention mechanism while underutilizing the role of the MLP and its non-linear components. As a result, our construction highlights how these non-linear MLP elements can play a critical role in enabling in-context learning, demonstrating their significance. We now present our main result.
@@
(Parameter count and bit precision) We note that the above construction utilizes $9(3+6S)^{2}+(S+T)(3+6S)$ parameters. A slightly more efficient construction in terms of parameter count with $8(3+6S)^{2}+(S+T)(3+6S)$ parameters can be found in App. B. For both these results, it suffices to have $\Omega(\log(T)+k)$ bits per parameter, with $\mathcal{O}(1/T)$ additive approximation error, where $T$ is the sequence length.
@@
We now provide an outline of the proof of Theorem 4, using notation consistent with the constructions in Theorems 3 and 3.
@@
In contrast to previous constructions, the MLP situated between the two attention layers plays a central role in our proof and is structured to reconstruct the complementary representation $\textbf{u}_{n}=\left({\sum_{i=0}^{k-1}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=0}^{k-1}3^{i}}\right)$, which corresponds to a hard attention pattern over the positions $\{n,n-1,\ldots,n-k+1\}$. Recall that this $\textbf{u}_{n}$ was constructed in 3 thanks to a second attention head, absent in our architecture.
@@
As in the second layer of Theorem 3, the attention mechanism in the second layer is configured to act as a $k$th-order induction head by computing an attention score at position $n$ that is proportional to the cosine similarity $\left({\langle\textbf{u}_{T},\textbf{v}_{n}\rangle/\|\textbf{u}_{T}\|_{2}\|\textbf{v}_{n}\|_{2}}\right)$. This score reaches its maximum value of one when $\textbf{u}_{T}=\textbf{v}_{n}$, and is strictly less than one otherwise. As the softmax temperature tends to infinity, or equivalently as the attention logits are scaled by a constant that tends to infinity, the resulting distribution converges to that of a conditional $k$-gram estimator. App. B details the complete proof. ∎
Importance of non-linearities. This construction demonstrates that non-linear components such as ReLU activations and layer normalization are not merely auxiliary, but play a critical role in enabling the transformer architecture to express higher-order inductive structures that are essential for in-context learning.
@@ 5 Gradient Descent Analysis
Thm. 4 establishes a representation result showing that a two-layer transformer with single attention heads can implement the conditional $k$-gram model. However, this does not guarantee that current optimization algorithms such as gradient descent (or its stochastic variants) can recover such a solution in practice. To this end, in this section we focus on first-order Markov chains and show that gradient descent on a reduced variant of our two-layer transformer indeed learns an induction head, and thereby the in-context conditional $1$-gram. On this note, we would like to emphasize that first-order Markov analysis a critical first-step and a building block for challenging higher-order analysis Nichani et al. (2024).
Building towards our result, we quickly recall a few recent works Edelman et al. (2024); Nichani et al. (2024) that also analyze the gradient dynamics on first-order Markov data with simplified transformer architectures. Specifically, Edelman et al. (2024) study an attention-only linear transformer without MLPs and softmax operations, while Nichani et al. (2024) analyze a disentangled transformer, with residual and attention streams split. On the other hand, our transformer architecture in Def. 2.2 consists of relative positional encodings, softmax attention, MLP with ReLU activations, and layer normalization. While non-linearities such as ReLU and layer normalization play a crucial role in our representation result, as illustrated in Sec. 4, at the same time, they make the gradient analysis challenging and intractable even for a simple first-order Markov setup.
To tackle this, we study a reduced model wherein we only treat the salient components of our two-layer architecture in Thm. 4, such as positional encodings and attention layer, as parameters, and treating the rest as approximately optimal. This enables for a a faithful reproduction of the ICL phenomenon exhibited by the parent model, whilst being tractable. This is akin to the reparameterization strategies of Makkuva et al. (2024b) and Nichani et al. (2024) for gradient-flow/gradient-descent analysis. Alternatively, this can also be viewed as a form of good parameter initialization, similar to that of Edelman et al. (2024); Nichani et al. (2024). We now start with our technical assumptions on the data distribution and the transformer architecture. We use notation from Sec. 2.2, dropping the head-identifying superscript.
@@
    Data distribution assumptions: Following Nichani et al. (2024), we assume the prior over transition kernels enforces non-degeneracy, positive transitions, and constant mean (see App. C).
@@
We minimize the cross-entropy loss for next-token prediction Nichani et al. (2024):
@@
Training algorithm. We denote the set of trainable parameters as $\theta=(\textbf{p},a_{2})$, where the positional scalars $\textbf{p}=[p_{0},p_{1},\ldots,p_{T}]^{T}$ and $a_{2}$ is the attention scalar. Similar to Nichani et al. (2024), we adopt a two-stage training procedure. (i) In the first stage, we optimize the positional scalars p using gradient descent for $T_{1}$ steps with learning rate $\eta_{1}$, (ii) in the second stage, we freeze p and train only $a_{2}$ for an additional $T_{2}$ steps using a separate learning rate $\eta_{2}$. A key distinction in our setup is how we treat layer normalization during training. While in the first stage we omit all layer norms in the MLP, during the second step we reintroduce them. This stage-wise enables a tractable theoretical analysis. We refer to App. C for further details. We now present our main result:
@@
Thus Thm. 5 illustrates that the transformer model trained via the two-stage algorithm achieves near-optimal loss asymptotically. Furthermore, we can also show that $\hat{\theta}$ approximates the conditional $1$-gram model, i.e. first-order induction head, with vanishing error as sequence length grows (see App. C). For completeness, perturbative analysis of the architecture consisting of two attention heads in the first layer and one in the second (see Sec. 3) for any-order Markov chains can be found in (App. D).
@@ Acknowledgements
We are grateful to Eshaan Nichani for insightful discussions and valuable input throughout the course of this project.
@@ A.1 Modification of the Layer Norm
Following Rajaraman et al. [2024] (Appendix C.1 in Rajaraman et al. [2024]), we use a modified version of layer normalization to extract the $\ell_{2}$-normalized direction of the embedding $\textbf{x}_{n}$. Specifically, we augment the embedding by concatenating it with its negation to form $\textbf{x}_{n};-\textbf{x}_{n}]$, which ensures that the mean of the features is zero. Applying standard layer normalization to this vector yields $\sqrt{d}\cdot\textbf{x}_{n}/\|\textbf{x}_{n}\|$ in the first $d$ components, since the standard deviation of the augmented vector is $\|\textbf{x}_{n}\|/\sqrt{d}$. To recover the unit-normalized embedding $\textbf{x}_{n}/\|\textbf{x}_{n}\|$, we simply divide the first $d$ components of the output by $\sqrt{d}$, and extract the first $d$ components. Throughout this work, we assume this normalization and scaling step is handled implicitly, and we interpret layer normalization as returning the normalized embedding direction. In order to avoid explicitly repeating this transformation at each layer, we redefine the layer norm to implicitly return $\textbf{v}/\|\textbf{v}\|_{2}$ for any input vector v. Hence, $\textbf{v}/\|\textbf{v}\|_{2}=\mathrm{LN}(\textbf{v})$, where, $\mathrm{LN}(\cdot)$ denotes the modified layer norm operator. Note that this modified definition is used throughout this paper.

### v2

@@ 3 Warm Up: Construction with Two Heads in Layer One, One in Layer Two
As a warm up, in this section, we prove our first main result that we can represent $k^{\text{th}}$-order Markov processes with a $2$-layer, 22-head transformer. Towards the same, we recall that the $k^{\text{th}}$-order in-context counting estimator, i.e. conditional $k$-gram model, defined in Subsec. 2.1 is closely related to the classical Laplacian smoothing, which in turn is the optimal Bayes estimator for the next-token predictive distribution [laplace1814essaiprob, rissanen1984]. In view of this fact, in order to estimate the optimal conditional distribution of each token in-context, it is natural to ask: how does a transformer represent this conditional $k$-gram model? To this end, rajaraman2024transformers introduces the notion of a $k$-th order induction head. This mechanism enables the attention layer of a transformer to assign at each time step the highest attention weight to past tokens whose length-$k$ context matches that of the current token. For the sake of completeness, we recall the formal definition.
@@
The salience of this higher-order induction head is immediately reflected by the fact that such a mechanism is indeed what is precisely needed to implement the conditional $k$-gram estimator from Subsec. 2.1. Capitalizing on this, rajaraman2024transformers proves the best known result for higher-order Markov processes, through a $3$-layer, 11-head transformer architecture, which we recall below to motivate our main result:
Theorem 1. (Theorem 4 in rajaraman2024transformers): The conditional $k$-gram model can be represented using a transformer with three layers, each containing a single attention head. The layers are separated by MLP blocks and include relative positional encodings as well as layer normalization. The embedding dimension scales as $\mathcal{O}(S)$.
@@
We provide a brief outline of the proof of Sec. 3. For each position $n$ in the sequence $s_{0:T}$, the first attention layer effectively learns a hard attention map concentrated on the indices $\{n,n-1,\ldots,n-k+1\}$. Specifically, within a subspace of its embedding space, this layer computes $\textbf{u}_{n}=\left({\sum_{i=0}^{k-1}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=0}^{k-1}3^{i}}\right)$, with attention weights $\operatorname{attn}_{n,i}=3^{i}/\left({\sum_{j=0}^{k-1}3^{j}}\right)$ for $i\in\{n,n-1,\ldots,n-k+1\}$, and zero otherwise. Intuitively, this $\textbf{u}_{n}$ allows to capture the past-$k$ context starting at position $n$. Similarly, the second attention layer focuses on the positions $\{n-1,n-2,\ldots,n-k\}$, and computes $\textbf{v}_{n}=\left({\sum_{i=1}^{k}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=1}^{k}3^{i}}\right)$, capturing the context shifted by one position. After these two layers, the embedding at position $n$ contains both $\textbf{v}_{n}/\|\textbf{v}_{n}\|_{2}$ and the normalized vector $\textbf{u}_{n}/\|\textbf{u}_{n}\|_{2}$, assuming appropriate choices of embedding dimensions, query/key/value matrices, positional encodings, and MLP layers with layer normalization. In particular, the architecture can be configured to compute the $\ell_{2}$-norm via layer norm, as detailed in Subsec. A.1. The third attention layer then functions as a $k$-th order induction head: at position $n$, the attention score is made proportional to the cosine similarity ${\langle\textbf{u}_{T},\textbf{v}_{n}\rangle}/\left({\|\textbf{u}_{T}\|_{2}\|\textbf{v}_{n}\|_{2}}\right)$. This quantity equals 1 when $\textbf{u}_{T}=\textbf{v}_{n}$, and is strictly less than 1 otherwise. As the softmax temperature tends to infinity, or equivalently, as the attention logits are scaled by a constant tending to infinity, the model approaches the behavior of a conditional $k$-gram estimator. Hence, the third layer effectively implements a $k$-th order induction head. ∎
@@
(Depth-width tradeoff) Sec. 3 thus reveals an interesting tradeoff between depth (number of attention layers) and width (maximum number of attention heads per layer). While our construction here has both depth and width two, its counterpart in Sec. 3 instead has depth three and width one. Thus our new architecture trades off depth with an additional attention head, thereby increasing the width.
@@
We provide a brief sketch of the proof for Sec. 3, using the same notation as that of the Sec. 3, and focusing on the key differences between the two architectures.
@@ 4 Main Result: Two-Layer Single-Head Construction
Figure 2: Attention maps learnt by a two-layer, single-head transformer trained on sequences generated by random Markov chains of order 33 (Subsec. 2.1). (i) in the first layer, the attention map shows a clear pattern: attention weights increase monotonically along the first three lower-diagonals and drop to zero beyond that. This suggests that the relative positional bias is maximized the diagonal with an offset of k=3-k=-3, i.e., the third diagonal below the main diagonal, which is consistent with the construction in Sec. 4, (ii) in the second layer, the attention map closely resembles the ideal attention pattern required to approximate the conditional $k$-gram estimator. We note that all experiments were conducted using standard initialization schemes. Additional experiments with different orders of markov chains and experimental details are provided in App. H. In Subsec. H.2, we also experimentally demonstrate that single-layer transformers fail to solve the induction head task with the same order of parameters. Finally, in Subsec. H.3, we test the robustness of the two-layer, single-head model to noise in the input sequences.
While Sec. 3 above demonstrates that the conditional $k$-gram can be represented by a two-layer transformer, it however relies on a two-head construction, thus still leaving open our motivating question: can a $2$-layer, 11-head transformer represent the conditional $k$-gram? In this section, we precisely address this gap and realize a $2$-layer, 11-head model to represent $k^{\text{th}}$-order Markov processes for any $k\geq 1$. The key idea behind our approach is to leverage the MLP—and specifically, the non-linearities such as ReLU and LayerNorm—to isolate symbols at particular positions that are useful for estimating induction heads. This contrasts with prior constructions, which primarily focused on the attention mechanism while underutilizing the role of the MLP and its non-linear components. As a result, our construction highlights how these non-linear MLP elements can play a critical role in enabling in-context learning, demonstrating their significance. We now present our main result.
@@
(Parameter count and bit precision) We note that the above construction utilizes $9(6S+3)^{2}+(6S+3)(2T+2S+9)$ parameters. A slightly more efficient construction in terms of parameter count with $8(6S+3)^{2}+(6S+3)(2T+2S+9)$ parameters can be found in App. B. For both these results, it suffices to have $\Omega(\log(T)+k)$ bits per parameter, with $\mathcal{O}(1/T)$ additive approximation error, where $T$ is the sequence length.
@@
We now provide an outline of the proof of Sec. 4, using notation consistent with the constructions in Sec. 3 and Sec. 3.
@@
In contrast to previous constructions, the MLP situated between the two attention layers plays a central role in our proof and is structured to reconstruct the complementary representation $\textbf{u}_{n}=\left({\sum_{i=0}^{k-1}3^{i}\cdot e^{S}_{x_{n-i}}}\right)/\left({\sum_{i=0}^{k-1}3^{i}}\right)$, which corresponds to a hard attention pattern over the positions $\{n,n-1,\ldots,n-k+1\}$. Recall that this $\textbf{u}_{n}$ was constructed in Sec. 3 thanks to a second attention head, absent in our architecture.
@@
As in the second layer of Sec. 3, the attention mechanism in the second layer is configured to act as a $k$th-order induction head by computing an attention score at position $n$ that is proportional to the cosine similarity $\left({\langle\textbf{u}_{T},\textbf{v}_{n}\rangle/\|\textbf{u}_{T}\|_{2}\|\textbf{v}_{n}\|_{2}}\right)$. This score reaches its maximum value of one when $\textbf{u}_{T}=\textbf{v}_{n}$, and is strictly less than one otherwise. As the softmax temperature tends to infinity, or equivalently as the attention logits are scaled by a constant that tends to infinity, the resulting distribution converges to that of a conditional $k$-gram estimator. App. B details the complete proof. ∎
Importance of non-linearities.
This construction demonstrates that non-linear components such as ReLU activations and layer normalization are not merely auxiliary, but play a critical role in enabling the transformer architecture to express higher-order inductive structures that are essential for in-context learning.
Note.
Due to the problem’s non-convexity, multiple constructions may enable two-layer, single-head transformers to implement a $k$th-order induction head; our approach offers one such construction motivated by two key insights. First, empirical observation: for datasets generated from $k$th-order Markov chains, we observe that in a trained transformer model, the first-layer attention weights increase monotonically along the first $k$ lower diagonals before dropping sharply to near-zero (see Fig. 2 and Subsec. H.1). Second, prior theoretical construction: as demonstrated in rajaraman2024transformers, the attention pattern can be designed to be proportional to a dyadic sum representation. Taken together, they offer both empirical and theoretical grounding for our construction of $k$th-order induction heads.
@@ 5 Gradient Descent Analysis
Sec. 4 establishes a representation result showing that a two-layer transformer with single attention heads can implement the conditional $k$-gram model. However, this does not guarantee that current optimization algorithms such as gradient descent (or its stochastic variants) can recover such a solution in practice. To this end, in this section we focus on first-order Markov chains and show that gradient descent on a reduced variant of our two-layer transformer indeed learns an induction head, and thereby the in-context conditional $1$-gram. On this note, we would like to emphasize that first-order Markov analysis a critical first-step and a building block for challenging higher-order analysis [nichani2024how].
Building towards our result, we quickly recall a few recent works [edelman2024evolution, nichani2024how] that also analyze the gradient dynamics on first-order Markov data with simplified transformer architectures. Specifically, edelman2024evolution study an attention-only linear transformer without MLPs and softmax operations, while nichani2024how analyze a disentangled transformer, with residual and attention streams split. On the other hand, our transformer architecture in Subsec. 2.2 consists of relative positional encodings, softmax attention, MLP with ReLU activations, and layer normalization. While non-linearities such as ReLU and layer normalization play a crucial role in our representation result, as illustrated in Sec. 4, at the same time, they make the gradient analysis challenging and intractable even for a simple first-order Markov setup.
To tackle this, we study a reduced model wherein we only treat the salient components of our two-layer architecture in Sec. 4, such as positional encodings and attention layer, as parameters, and treating the rest as approximately optimal. This enables for a a faithful reproduction of the ICL phenomenon exhibited by the parent model, whilst being tractable. This is akin to the reparameterization strategies of makkuva2024local and nichani2024how for gradient-flow/gradient-descent analysis. Alternatively, this can also be viewed as a form of good parameter initialization, similar to that of [edelman2024evolution, nichani2024how]. We now start with our technical assumptions on the data distribution and the transformer architecture. We use notation from Subsec. 2.2, dropping the head-identifying superscript.
@@
    Data distribution assumptions: Following [nichani2024how], we assume the prior over transition kernels enforces non-degeneracy, positive transitions, and constant mean (see App. C).
@@
We minimize the cross-entropy loss for next-token prediction [nichani2024how]:
@@
Training algorithm. We denote the set of trainable parameters as $\theta=(\textbf{p},a_{2})$, where the positional scalars $\textbf{p}=[p_{0},p_{1},\ldots,p_{T}]^{T}$ and $a_{2}$ is the attention scalar. Similar to [nichani2024how], we adopt a two-stage training procedure. (i) In the first stage, we optimize the positional scalars p using gradient descent for $T_{1}$ steps with learning rate $\eta_{1}$, (ii) in the second stage, we freeze p and train only $a_{2}$ for an additional $T_{2}$ steps using a separate learning rate $\eta_{2}$. A key distinction in our setup is how we treat layer normalization during training. While in the first stage we omit all layer norms in the MLP, during the second step we reintroduce them. This stage-wise enables a tractable theoretical analysis. We refer to App. C for further details. We now present our main result:
@@
Thus Sec. 5 illustrates that the transformer model trained via the two-stage algorithm achieves near-optimal loss asymptotically. Furthermore, we can also show that $\hat{\theta}$ approximates the conditional $1$-gram model, i.e. first-order induction head, with vanishing error as sequence length grows (see App. C). For completeness, perturbative analysis of the architecture consisting of two attention heads in the first layer and one in the second (see Sec. 3) for any-order Markov chains can be found in App. D.
@@ Acknowledgements
This work was supported in part by the Swiss National Science Foundation under Grant No. 200364. We are grateful to Eshaan Nichani for insightful discussions and valuable input throughout the preparation of this paper.
@@ A.1 Modification of the Layer Norm
Following rajaraman2024transformers (Appendix C.1 in rajaraman2024transformers), we use a modified version of layer normalization to extract the $\ell_{2}$-normalized direction of the embedding $\textbf{x}_{n}$. Specifically, we augment the embedding by concatenating it with its negation to form $\textbf{x}_{n};-\textbf{x}_{n}]$, which ensures that the mean of the features is zero. Applying standard layer normalization to this vector yields $\sqrt{d}\cdot\textbf{x}_{n}/\|\textbf{x}_{n}\|$ in the first $d$ components, since the standard deviation of the augmented vector is $\|\textbf{x}_{n}\|/\sqrt{d}$. To recover the unit-normalized embedding $\textbf{x}_{n}/\|\textbf{x}_{n}\|$, we simply divide the first $d$ components of the output by $\sqrt{d}$, and extract the first $d$ components. Throughout this work, we assume this normalization and scaling step is handled implicitly, and we interpret layer normalization as returning the normalized embedding direction. In order to avoid explicitly repeating this transformation at each layer, we redefine the layer norm to implicitly return $\textbf{v}/\|\textbf{v}\|_{2}$ for any input vector v. Hence, $\textbf{v}/\|\textbf{v}\|_{2}=\mathrm{LN}(\textbf{v})$, where, $\mathrm{LN}(\cdot)$ denotes the modified layer norm operator. Note that this modified definition is used throughout this paper.